# The Latency Price of Threshold Cryptosystem in Blockchains

### v1

@@ 5 Blockchain-Native Threshold Cryptosystem with Tight Thresholds
1:upon ${\sf prefinalize}(r,m)$ do
@@
3:    send $({\tt PREFIN},r,m)$ and $({\tt SHARE},r,m,\sigma_{i})$ to all parties
4:upon receiving $({\tt PREFIN},r,m)$ from party $j$ do
5:    $T_{r,m}\leftarrow T_{r,m}\cup\{j\}$
6:    if $|T_{r,m}|\geq t_{\sf fin}$ and $\bm{m}[r]=\{\}$ then
7:         call ${\sf finalize}(r,m)$
8:upon ${\sf finalize}(r,m)$ do
9:    let $\bm{m}[r]\leftarrow m$
10:    $queue.push(r)$
11:    if $({\tt SHARE},r,*,*)$ not sent then
12:         let $\sigma_{i}\leftarrow{\sf PEval}(\llbracket{s}\rrbracket_{i},(r,m))$
13:         send $({\tt SHARE},r,m,\sigma_{i})$ to each other party
@@
In this section, we present a protocol for $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ (Definition 13) for tight thresholds that has low latency. In any given round $r\in\mathbb{N}$ and in any execution, the global finalization time of our protocol is same as the global reconstruction time, i.e., ${\sf GFT}_{r}={\sf GRT}_{r}$. Moreover, in error-free executions, honest parties in our protocol learns the $\sf TC$ output simultaneously with the $\sf MBB_{FT}$ output, i.e., ${\sf L}_{r}=0$. We summarize our construction in Algorithm 2 and describe it next.
5.1 Design
@@
The setup phase is identical to that of Algorithm 1, except that the secrecy and reconstruction thresholds are set to be equal to the finalization threshold of $\sf MBB_{FT}$. Note that, the Termination property of $\sf MBB_{FT}$ requires $t_{\sf fin}\leq n-t$, as honest parties needs to finalize a message even when corrupted parties do not send any throughout the protocol. Next, unlike Algorithm 1, parties reveal their $\sf TC$ output shares when they prefinalize a message. More specifically, for every round $r$, each party $i$ computes $\sigma_{i}:={\sf PEval}(\llbracket{s}\rrbracket_{i},(r,m))$ upon prefinalizing the value $(r,m)$, and sends the SHARE message $({\tt SHARE},r,m,\sigma_{i})$ to all parties in addition to sending the PREFIN message $({\tt PREFIN},r,m)$. When a party receives $t_{\sf fin}$ PREFIN messages $({\tt PREFIN},r,m)$, it finalizes the message $m$ for round $r$, by adding round $r$ to $queue$ and recording $m$ in $\bm{m}[r]$. The party also computes and sends $\sigma_{i}$ if it has not done so. The reconstruction and output phases are also identical to Algorithm 1, where parties collect and combine shares to generate $\sf TC$ output, and output the result round-by-round.
5.2 Analysis of Algorithm 2
Theorem 5.1
Proof
Let $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ denote the protocol.
Secrecy.
We first prove that, for any round $r\in\mathbb{N}$, if an honest party outputs $(r,m,\sigma)$ then $\sigma={\sf Eval}(s,(r,m))$. For the sake of contradiction, suppose that for some $r\in\mathbb{N}$, an honest party outputs $(r,m,\sigma^{\prime})$ where $\sigma^{\prime}\neq{\sf Eval}(s,(r,m))$. By the protocol, the party outputs ${\sf finalize}(r,m)$ in $\sf MBB_{FT}$. Also, according to the protocol and the Robustness property of $\sf TC$, $\sigma^{\prime}={\sf Eval}(s,(r,m^{\prime}))$ for some $m^{\prime}\neq m$. This implies, by the Unpredictability property of $\sf TC$, that at least $t_{\sf sec}-|\mathcal{C}|$ honest parties have revealed their $\sf TC$ output shares ${\sf PEval}(\llbracket{s}\rrbracket,(r,m^{\prime}))$. Let $T_{m^{\prime}}$ be set indicating these honest parties. Note that in Algorithm 2 an honest party reveals its $\sf TC$ output share for any $(r,m^{\prime})$ only upon prefinalizing $(r,m^{\prime})$ or finalizing $(r,m^{\prime})$. Consider the latter case: if any party $i\in T_{m^{\prime}}$ finalizes a message $(r,m^{\prime})$ for $m^{\prime}\neq m$, then this violates the Agreement property of $\sf MBB_{FT}$, and hence a contradiction. This implies that all parties in $T_{m^{\prime}}$ has called ${\sf prefinalize}(r,m^{\prime})$ in $\sf MBB_{FT}$. However, since $\sf MBB_{FT}$ has the finalization threshold $t_{\sf fin}$, $t_{\sf fin}=t_{\sf sec}$ and $|T_{m^{\prime}}|\geq t_{\sf sec}-|\mathcal{C}|=t_{\sf fin}-|\mathcal{C}|$ implies that the message $m^{\prime}$ is globally finalized. This again violates the Agreement property of $\sf MBB_{FT}$, hence a contradiction. Therefore, for any round $r\in\mathbb{N}$, if an honest party outputs $(r,m,\sigma)$ then $\sigma={\sf Eval}(s,(r,m))$.
Agreement.
Termination.
Validity.
Total Order.
Theorem 5.2
Proof
Recall from Section 3, for any round $r$, ${\sf GFT}_{r}$ and ${\sf GRT}_{r}$ are the global finalization time and global reconstruction time of round $r$, respectively. Consider any execution. By the definition of ${\sf GFT}_{r}$, $t_{\sf fin}-|\mathcal{C}|$ honest parties have called ${\sf prefinalize}(r,m)$ at time ${\sf GFT}_{r}$. According to Algorithm 2, these honest parties have also revealed their $\sf TC$ output shares for $(r,m)$ at time ${\sf GFT}_{r}$. Since, $t_{\sf rec}=t_{\sf fin}$, this implies that the adversary can adversary to reconstruct the $\sf TC$ output at time ${\sf GFT}_{r}$, and hence ${\sf GRT}_{r}\leq{\sf GFT}_{r}$. From Lemma 1, we have that ${\sf GRT}_{r}\geq{\sf GFT}_{r}$. Therefore, we conclude that ${\sf GFT}_{r}={\sf GRT}_{r}$ for any $r$ in any execution.
Theorem 5.3
Proof
6 Blockchain-Native Threshold Cryptography with Ramp Thresholds
In this section, we present the results for $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ (Definition 13) with ramp thresholds $t_{\sf sec}<t_{\sf rec}$. We first show in Section 6.1 an impossibility result that, for any round $r$, for any $\sf BTC_{FT}$ with ramp thresholds, there must exist an execution where ${\sf GRT}_{r}>{\sf GFT}_{r}$. To circumvent the impossibility result, in Section 6.2 we propose an optimistic protocol that achieves ${\sf GRT}_{r}={\sf GFT}_{r}$ in synchronous executions and ${\sf L}_{r}=0$ in optimistic executions.
@@ 6.1 Impossibility
Proof
For the sake of contradiction, suppose that there exists a protocol $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ with ramp thresholds, such that for any given round $r$, in all executions we have ${\sf GRT}_{r}={\sf GFT}_{r}$. Let $\mathcal{E}$ be one such execution and let $\tau_{\mathcal{E}}={\sf GRT}_{r}={\sf GFT}_{r}$. Let $s$ denote the secret shared between the parties. Since $\sf MBB_{FT}$ has the finalization threshold $t_{\sf fin}$, at least $t_{\sf fin}-|\mathcal{C}|$ honest parties have called ${\sf prefinalize}(r,m)$ at $\tau_{\mathcal{E}}$. Similarly, by the definition of ${\sf GRT}_{r}$, at least $t_{\sf rec}-|\mathcal{C}|$ honest parties have revealed their $\sf TC$ output shares ${\sf PEval}(\llbracket{s}\rrbracket,(r,m))$ at $\tau_{\mathcal{E}}$. Without loss of generality, we can assume that exactly $t_{\sf fin}-|\mathcal{C}|$ honest parties have called ${\sf prefinalize}(r,m)$ and $t_{\sf rec}-|\mathcal{C}|$ honest parties have revealed their $\sf TC$ output shares at $\tau_{\mathcal{E}}$. Let $h$ be any honest party that called ${\sf prefinalize}(r,m)$ at time $\tau_{\mathcal{E}}$ in the execution $\mathcal{E}$. There are two cases, which we denote with (1) and (2) below.
  1. (1)
    $h$ also reveals its $\sf TC$ output share ${\sf PEval}(\llbracket{s}\rrbracket_{h},(r,m))$ at time $\tau_{\mathcal{E}}$. Consider another execution $\mathcal{E}^{\prime}$ identical to $\mathcal{E}$ up to time $\tau_{\mathcal{E}}$ with the only difference that, $h$ calls ${\sf prefinalize}(r,m)$ and reveals ${\sf PEval}(\llbracket{s}\rrbracket_{h},(r,m))$ at time time $\tau_{\mathcal{E}}+\epsilon$ for some $\epsilon>0$ (say due to asynchrony in computation or communication). Therefore, $\tau_{\mathcal{E}}+\epsilon$ is the global finalization time of execution $\mathcal{E}^{\prime}$.
    Now, consider the time $\tau_{\mathcal{E}}$ in the new execution $\mathcal{E}^{\prime}$. The message $m$ is not globally finalized at $\tau_{\mathcal{E}}$ for $r$ since only $t_{\sf fin}-|\mathcal{C}|-1$ honest parties have called ${\sf prefinalize}(r,m)$. However, since $t_{\sf rec}>t_{\sf sec}$, there are $t_{\sf rec}-|\mathcal{C}|-1\geq t_{\sf sec}-|\mathcal{C}|$ honest parties that have revealed their $\sf TC$ output shares ${\sf PEval}(\llbracket{s}\rrbracket,(r,m))$. This implies that in execution $\mathcal{E}^{\prime}$ the adversary $\mathcal{A}$ can compute the $\sf TC$ output ${\sf Eval}(s,(m,r))$ at $\tau_{\mathcal{E}}$, which is earlier than its global finalization time $\tau_{\mathcal{E}}+\epsilon$. This contradicts Lemma 1.
  2. (2)
    $h$ does not reveal its $\sf TC$ output share at time $\tau_{\mathcal{E}}$. Similarly, consider another execution $\mathcal{E}^{\prime}$ identical to $\mathcal{E}$ up to time $\tau_{\mathcal{E}}$ with the only difference that, $h$ calls ${\sf prefinalize}(r,m)$ at time $\tau_{\mathcal{E}}+\epsilon$ for some $\epsilon>0$ (say due to asynchrony in computation or communication). Therefore, for the new execution $\mathcal{E}^{\prime}$, $\tau_{\mathcal{E}}+\epsilon$ is the global finalization time, and the global reconstruction time remains $\tau_{\mathcal{E}}$. Then, for the new execution $\mathcal{E}^{\prime}$, the global reconstruction comes before the global finalization, contradicting Lemma 1. ∎
Theorem 6.1 states that for any blockchain-native threshold cryptosystem with finalization threshold and ramp thresholds, there always exists an execution where global reconstruction occurs after global finalization for each round. In fact, existing solutions for $\sf BTC$ with ramp thresholds all have a latency of at least one message delay, such as Das et al. [23].
@@ 6.2 Fast Path
Theorem 6.1 claims that no $\sf BTC_{FT}$ protocol with ramp thresholds can always guarantee that the global finalization and reconstruction occur simultaneously, implying that the latency of $\sf BTC_{FT}$ may be unavoidable. Fortunately, we can circumvent this impossibility result in optimistic executions. In this section, we describe a simple protocol named fast path that, for any round $r$, achieves ${\sf GRT}_{r}={\sf GFT}_{r}$ under synchronous executions (Definition 15), and ${\sf L}_{r}=0$ under optimistic executions (Definition 16). As we illustrate in Section 7.3, in practice, our new protocol achieves significantly lower latency compared to the strawman protocol.
[…]
We address this issue as follows: First, we share the $\sf TC$ secret $s$ among the parties twice, using independent randomness, with two different pairs of thresholds $(t_{\sf sec},t_{\sf rec})$ and $(t_{\sf rec}^{\prime},t_{\sf sec}^{\prime})$. In particular, we choose the thresholds $(t_{\sf sec},t_{\sf rec})$ with the constraint that $t_{\sf sec}\geq t+1$ and $t_{\sf rec}\leq n-t$. Similarly, we choose $(t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})$ with the constraint that $t_{\sf sec}^{\prime}\geq t_{\sf fin}$ and $t_{\sf rec}^{\prime}\leq n$. Let $\{\llbracket{s}\rrbracket\}_{i\in[n]}$ and $\{\llbracket{s}\rrbracket^{\prime}\}_{i\in[n]}$ be the secret shares of $s$ with thresholds $(t_{\sf sec},t_{\sf rec})$ and $(t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})$, respectively. Second, we add an fast path, where parties reveal their $\sf TC$ output shares they compute with $\{\llbracket{s}\rrbracket^{\prime}\}_{i\in[n]}$ immediately upon prefinalizing a message.
We present our final protocol in Algorithm 3 and discuss it next. The setup phase is similar to Algorithm 1, except that the same secret $s$ is shared twice, using independent randomness, for the slow path and fast path, respectively. For any round $r$, each party does the following:
  • Fast path: For the fast path, when a party prefinalizes a message $m$ for round $r$, it reveals its $\sf TC$ output share ${\sf PEval}(\llbracket{s}\rrbracket^{\prime}_{i},(r,m))$. Once a party receives $t_{\sf rec}^{\prime}$ verified shares of the fast path, it reconstructs the $\sf TC$ output.
  • Slow path: For the slow path, upon $\sf MBB_{FT}$ finalization for message $m$ and round $r$, each party $i$ reveals its $\sf TC$ output share ${\sf PEval}(\llbracket{s}\rrbracket_{i},(r,m))$. Next, any party who has not received $t_{\sf rec}^{\prime}$ verified $\sf TC$ output shares from the fast path waits to receive $t_{\sf rec}$ verified $\sf TC$ output shares from the slow path. Once the party receives $t_{\sf rec}$ verified shares of the slow path, it reconstructs the $\sf TC$ output.
Lastly, similarly to the output phase of Algorithm 1, to guarantee Total Order, the parties push the finalized rounds into the First-in-first-out (FIFO) $queue$ and output the result round-by-round once either the fast path or slow path has reconstructed the $\sf TC$ output. So the latency of the protocol is the minimum latency of the two paths.
Note that, a party reveals its share of the slow path even if it has already revealed its share of the fast path or reconstructed the $\sf TC$ output from the fast path. This is crucial for ensuring Termination, because with corrupted parties sending their $\sf TC$ output shares to only a subset $\mathcal{S}$ of honest parties, it is possible that only parties in $\mathcal{S}$ can reconstruct the $\sf TC$ output from the fast path. If the honest parties in $\mathcal{S}$ do not reveal their $\sf TC$ output shares of the slow path, the remaining honest parties would be unable to reconstruct the $\sf TC$ output, thereby losing the Termination guarantee.
Algorithm 3 Fast Path for Blockchain-Native Threshold Cryptosystem with Ramp Thresholds
@@
3:let ($(\{{\sf pk}_{j}\}_{j\in[n]},\llbracket{s}\rrbracket_{i})\leftarrow{\sf ShareGen}(s)$ for thresholds $t+1\leq t_{\sf sec}<t_{\sf rec}\leq n-t$triangleright For slow path
4:let ($(\{{\sf pk}^{\prime}_{j}\}_{j\in[n]},\llbracket{s}\rrbracket^{\prime}_{i})\leftarrow{\sf ShareGen}(s)$ for thresholds $t_{\sf sec}^{\prime}=t_{\sf fin}<t_{\sf rec}^{\prime}\leq n$triangleright For fast path
@@
1:upon ${\sf prefinalize}(r,m)$ do
@@
1:upon receiving $({\tt PREFIN},r,m)$ from party $j$ do
2:    $T_{r,m}\leftarrow T_{r,m}\cup\{j\}$
3:    if $|T_{r,m}|\geq t_{\sf fin}$ and $\bm{m}[r]=\{\}$ then
4:         call ${\sf finalize}(r,m)$
5:upon ${\sf finalize}(r,m)$ do
6:    let $\bm{m}[r]\leftarrow m$
7:    $queue.push(r)$
@@
6.3 Proof of Correctness
Theorem 6.2

### v3

@@ 4 Tight Blockchain-Native Threshold Cryptosystem
1:upon sending $({\tt PREFIN},r,m)$ to all parties do $\triangleright$ Augment the consensus protocol
@@
3:    send $({\tt SHARE},r,m,\sigma_{i})$ to all parties
4:upon ${\sf finalize}(r,m)$ do
5:    let $\bm{m}[r]\leftarrow m$
6:    $queue.push(r)$
7:    if $({\tt SHARE},r,*,*)$ not sent then
9:         send $({\tt SHARE},r,m,\sigma_{i})$ to each other party
@@
In this section, we present a protocol for $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ (Definition 10) for tight thresholds that has low latency. In any round $r\in\mathbb{N}$ of any execution, the global finalization time of our protocol is same as the global reconstruction time, i.e., ${\sf GFT}_{r}={\sf GRT}_{r}$. Moreover, in error-free executions  [^3] We require error-free for the ${\sf L}_{r}=0$ claim, otherwise malicious parties may cause honest parties to prefinalize at different times and lead to ${\sf L}_{r}>0$., honest parties in our protocol learns the $\sf TC$ output simultaneously with the $\sf MBB_{FT}$ output, i.e., ${\sf L}_{r}=0$. We summarize our construction in Algorithm 2 and describe it next.
@@
The setup phase is identical to that of Algorithm 1, except that the secrecy and reconstruction thresholds are set to be equal to the finalization threshold of $\sf MBB_{FT}$. Note that, the Termination property of $\sf MBB_{FT}$ requires $t_{\sf fin}\leq n-t$, as honest parties needs to finalize a message even when corrupted parties do not send any throughout the protocol. Next, unlike Algorithm 1, parties reveal their $\sf TC$ output shares when they prefinalize a message. More specifically, for every round $r$, each party $i$ computes $\sigma_{i}\leftarrow{\sf PEval}(\llbracket{s}\rrbracket_{i},(r,m))$ upon prefinalizing the value $(r,m)$, and sends the SHARE message $({\tt SHARE},r,m,\sigma_{i})$ to all parties in addition to sending the PREFIN message $({\tt PREFIN},r,m)$. When a party receives $t_{\sf fin}$ PREFIN messages $({\tt PREFIN},r,m)$, it finalizes the message $m$ for round $r$, by adding round $r$ to $queue$ and recording $m$ in $\bm{m}[r]$. The party also computes and sends $\sigma_{i}$ if it has not done so. The reconstruction and output phases are also identical to Algorithm 1, where parties collect and combine shares to generate $\sf TC$ output, and output the result round-by-round. We defer the protocol analysis to Appendix 0.C due to space constraints.
5 Ramp Blockchain-Native Threshold Cryptography
In this section, we present an impossibility result and a feasibility result for $\sf BTC_{FT}=(\sf MBB_{FT},\sf TC)$ (Definition 10) with ramp thresholds $t_{\sf sec}<t_{\sf rec}$.
@@ 5.1 Impossibility
Due to space constraints, the proof of Theorem 5.1 is deferred to Section 0.D.1. Theorem 5.1 states that for any blockchain-native threshold cryptosystem with finalization threshold and ramp thresholds, there always exists an execution where global reconstruction occurs after global finalization for each round. In fact, existing solutions for $\sf BTC$ with ramp thresholds all have a latency of at least one message delay, such as Das et al. [23].
@@ 5.2 Fast Path
Algorithm 3 Fast Path for Ramp Blockchain-Native Threshold Cryptosystem
@@
1:let $\bm{m}\leftarrow\{\}$, $\bm{\sigma}\leftarrow\{\}$, $queue\leftarrow\{\}$
2:let ($(\{{\sf pk}_{j}\}_{j\in[n]},\llbracket{s}\rrbracket_{i})\leftarrow{\sf ShareGen}(s)$ for $t+1\leq t_{\sf sec}<t_{\sf rec}\leq n-t$triangleright For slow path
3:let ($(\{{\sf pk}^{\prime}_{j}\}_{j\in[n]},\llbracket{s}\rrbracket^{\prime}_{i})\leftarrow{\sf ShareGen}(s)$ for $t_{\sf sec}^{\prime}=t_{\sf fin},t_{\sf rec}^{\prime}=\min(t_{\sf sec}^{\prime}+(t_{\sf rec}-t_{\sf sec}),n)$triangleright For fast path, $s$ is unknown to any party
@@
1:upon sending $({\tt PREFIN},r,m)$ to all parties do $\triangleright$ Augment the consensus protocol
@@
1:upon ${\sf finalize}(r,m)$ do
2:    let $\bm{m}[r]\leftarrow m$ and $\sigma_{i}\leftarrow{\sf PEval}(\llbracket{s}\rrbracket_{i},(r,m))$
3:    $queue.push(r)$
@@
Theorem 5.1 claims that no $\sf BTC_{FT}$ protocol with ramp thresholds can always guarantee that the global finalization and reconstruction occur simultaneously, implying that the latency of $\sf BTC_{FT}$ may be unavoidable. Fortunately, we can circumvent this impossibility result in optimistic executions. In this section, we describe a simple protocol named fast path that, for any round $r$, achieves ${\sf GRT}_{r}={\sf GFT}_{r}$ under synchronized executions, and ${\sf L}_{r}=0$ under optimistic executions  [^4] Similar to Section 4, we require error-free for the ${\sf L}_{r}=0$ claim. However, the honest parties can reconstruct the $\sf TC$ output via fast path as long as $|{\mathcal{H}}|\geq t_{\sf rec}^{\prime}$ (with ${\sf L}_{r}>0$).. As we illustrate in Section 6.1, in practice, our new protocol achieves significantly lower latency compared to the strawman protocol.
The key observation from Theorem 5.1 is that, to ensure the same global finalization and reconstruction time, the secrecy threshold cannot be lower than the finalization threshold; otherwise, the $\sf TC$ output could be revealed before $\sf MBB_{FT}$ finalizes a message. A naive way to address this is to increase the secrecy threshold to match the finalization threshold, i.e., $t_{\sf sec}=t_{\sf fin}$. However, the issue is that, since the threshold is ramped, $t_{\sf rec}>t_{\sf sec}=t_{\sf fin}=n-t$ ($\sf MBB_{FT}$ protocols with optimal resilience typically have $t_{\sf fin}=n-t$ to ensure quorum intersection for safety), there may not be enough honest shares for parties to reconstruct $\sf TC$ output upon finalizing the $\sf MBB_{FT}$ output, violating the Termination property.
We address this issue as follows: First, we share the $\sf TC$ secret $s$ among the parties twice, using independent randomness, with two different pairs of thresholds $(t_{\sf sec},t_{\sf rec})$ and $(t_{\sf rec}^{\prime},t_{\sf sec}^{\prime})$. The new thresholds are set to be $t_{\sf sec}^{\prime}=t_{\sf fin},t_{\sf rec}^{\prime}=\min(t_{\sf sec}^{\prime}+(t_{\sf rec}-t_{\sf sec}),n)$. Let $\{\llbracket{s}\rrbracket\}_{i\in[n]}$ and $\{\llbracket{s}\rrbracket^{\prime}\}_{i\in[n]}$ be the secret shares of $s$ with thresholds $(t_{\sf sec},t_{\sf rec})$ and $(t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})$, respectively. Second, we add a fast path, where parties reveal their $\sf TC$ output shares they compute with $\{\llbracket{s}\rrbracket^{\prime}\}_{i\in[n]}$ immediately upon prefinalizing a message.
Our final protocol is in Algorithm 3. The setup phase is similar to Algorithm 1, except the same secret $s$ is shared twice, using independent randomness, for the slow path and fast path, respectively. Each party does the following:
  • Fast path: When a party prefinalizes a message $m$ for round $r$, it reveals its $\sf TC$ output share ${\sf PEval}(\llbracket{s}\rrbracket^{\prime}_{i},(r,m))$. Once a party receives $t_{\sf rec}^{\prime}$ verified shares of the fast path, it reconstructs the $\sf TC$ output.
  • Slow path: Upon $\sf MBB_{FT}$ finalization for message $m$ and round $r$, each party $i$ reveals its $\sf TC$ output share ${\sf PEval}(\llbracket{s}\rrbracket_{i},(r,m))$. Next, any party who has not received $t_{\sf rec}^{\prime}$ verified $\sf TC$ output shares from the fast path waits to receive $t_{\sf rec}$ verified $\sf TC$ output shares from the slow path. Once the party receives $t_{\sf rec}$ verified shares of the slow path, it reconstructs the $\sf TC$ output.
Lastly, similarly to Algorithm 1, to guarantee Total Order, the parties push the finalized rounds into the FIFO $queue$ and output the result round-by-round once either the fast path or slow path has reconstructed the $\sf TC$ output. So the latency of the protocol is the minimum latency of the two paths.
Note that, a party reveals its share of the slow path even if it has revealed its share of the fast path or reconstructed the $\sf TC$ output from the fast path. This is crucial for ensuring Termination, because with corrupted parties sending their $\sf TC$ output shares to only a subset $\mathcal{S}$ of honest parties, it is possible that only parties in $\mathcal{S}$ can reconstruct the $\sf TC$ output from the fast path. If honest parties in $\mathcal{S}$ do not reveal their shares of the slow path, the remaining honest parties cannot reconstruct the $\sf TC$ output, thereby losing the Termination guarantee. We defer the protocol analysis to Appendix 0.E due to space constraints.
6 Distributed Randomness: A Case Study
In this section, we implement and evaluate distributed randomness as a concrete example of blockchain-native threshold cryptosystem, to demonstrate the effectiveness of our solution in reducing the latency for real-world blockchains. We implement the fast path (Algorithm 3) for Das et al. [23], which is a distributed randomness scheme designed for proof-of-stake blockchains and is deployed in the Aptos blockchain [11] to enable smart contracts to use randomness [9]. We then compare our latency (using both micro-benchmarks and end-to-end evaluation) with the [23], that implements the strawman protocol (Algorithm 1).
In the rest of the section, we first provide a very brief overview of [23] and our implementation of fast-path protocol atop their scheme. Due to space constraints, more details are deferred to Appendix 0.E.
Overview of Das et al. [23]. Das et al. [23] is a distributed randomness protocol for proof-of-stake blockchains where each party has a (possibly unequal) stake, and the blockchain is secure as long as the adversary corrupts parties with combined stake less than $1/3$-th of the total stake. Since the total stake in practice can be very large, [23] first assigns approximate stakes of parties to a much smaller value called weights, and this process is called rounding. Parties in [23] then participate in a publicly verifiable secret sharing (PVSS) based distributed key generation (DKG) protocol to receive secret shares of a $\sf TC$ secret $s$. After DKG, Das et al. [23] implements the weighted extension of slow path (Algorithm 1) for blockchain-native threshold cryptosystem (Definition 10), where they use a distributed verifiable unpredictable function (VUF) as the $\sf TC$ protocol. More precisely, for each finalized block, each party computes and reveals its VUF shares. Next, once a party receives verified VUF shares from parties with combined weights greater than or equal to $w$, it reconstructs the VUF output.
Implementation of fast path (Algorithm 3). Recall that the fast path requires sharing the same secret with two sets of thresholds, i.e., $t_{\sf sec}<t_{\sf rec}$ for the slow path and $t_{\sf sec}^{\prime}<t_{\sf rec}^{\prime}$ for the fast path. Consequently, we augment the rounding algorithm of [23] to additionally take $(t_{\sf sec}^{\prime},t_{\sf rec}^{\prime})$ as input, and output the weight threshold for the fast path. To setup the secret-shares of the $\sf TC$ secret, we use the DKG protocol of [23] with the following minor modifications. Each party starts by sharing the same secret independently using two weight thresholds $w$ and $w^{\prime}$. The rest of the DKG protocol is identical to [23], except parties agree on two different aggregated PVSS transcript instead of one. Note that, these doubles the computation and communication cost of DKG. As described in Algorithm 3, parties reveal their VUF shares ($\sf TC$ output shares) for the fast path upon prefinalizing a block, and for the slow path upon finalizing a block. For both paths, the parties collect the VUF shares and are ready to execute the block as soon as the randomness ($\sf TC$ output) is reconstructed from either path.
6.1 Evaluation Results.
Scheme Randomness Latency Consensus Latency Overhead on top of Consensus
Das et al. [23] 85.5 ms 362 ms 23.6%
our fast-path 24.7 ms 362 ms 6.8%
Table 2: Latencies of Das et al. [23] and our fast-path on Aptos mainnet (July 2024).
We implement our fast-path protocol (Algorithm 3) in Rust, atop the open-source Das et al. [23] implementation [11] on the Aptos blockchain. We worked with Aptos Labs to deploy and evaluate our protocol on the Aptos mainnet.
Setup and metrics. As of July 2024, the Aptos blockchain is run by $140$ validators, distributed $50$ cities across $22$ countries with the stake distributed described in [4]. The $50$-th, $70$-th and $90$-th percentile (average) of round-trip latency between the blockchain validators is approximately $150$ms, $230$ms, and $400$ms, respectively. Due to space limitation, we defer other details of the evaluation setup to Section 0.E.3. We measure the randomness latency as the duration required to generate randomness for each block, as in Definition 9. It measures the duration from the moment the block is finalized by consensus to the when the randomness for that block becomes available. We report the average randomness latency (measured over a period of 12 hours). We also measure and compare the setup overhead for Das et al. [23] and fast-path, using micro-benchmarks on machines of the same hardware specs as the Aptos mainnet.
Randomness latency. Table 2 summarizes the latency comparison of our fast-path and Das et al. [23]. As observed, fast-path significantly reduces the randomness latency of Das et al. by 71%, from 85.5 ms to 24.7 ms. As mentioned in Section 5.2, the small latency overhead of fast-path comes from the fact that honest parties may need to wait slightly longer after local finalization to receive additional shares from the fast path, since the reconstruction threshold of the fast path is higher than the finalization threshold. To show the significance of the latency improvement for consensus end-to-end latency, we also measure the consensus latency 362 ms as the duration of each block from proposed to finalized. As shown in the table, fast-path improves the latency overhead from $23.6\%$ to $6.8\%$ in terms of the consensus latency. The latency of the slow path in our fast-path protocol (Algorithm 3) is comparable to that of Das et al. [23] and is therefore omitted for brevity.
7 Related Work
Threshold cryptography in blockchains. In many applications of modern blockchains, threshold cryptography plays a vital role. An important example of blockchain-native threshold cryptosystem is to generate distributed randomness for blockchains [23, 31]. where existing real-world deployments [23, 31, 11, 48, 28] follow the threshold VRF-based approach. Numerous blockchain research
[40, 50, 13, 52, 39, 33, 37, 17] focuses on MEV countermeasures and privacy enhancement using threshold decryption, where transactions are encrypted with threshold decryption and only revealed and executed by blockchain parties after finalization. To the best of our knowledge, all existing constructions supporting the blockchain-native threshold cryptosystem mentioned above incur an additional round of latency, except [17]. In [17], threshold decryption is employed to mitigate MEV while using Tendermint as the consensus protocol, with parties revealing their decryption shares when voting for a block. However, the described approach is insecure, as an adversary could decrypt the block before it is finalized. This vulnerability arises from an incorrect description of the Tendermint consensus mechanism, which mistakenly assumes only a single round of consensus voting. Furthermore, their approach is less general compared to the results presented in this paper.
Acknowledgments
Appendix 0.A Preliminaries
0.A.1 Cryptography Definitions
[…]
Game 0.A.1 (Robustness Under Chosen Message Attack)
[…]
    1. 1.
    2. 2.
Definition 11 (Robustness Under Chosen Message Attack)
[…]
Next, we describe the secrecy property. Intuitively, the secrecy property ensures that ${\sf Eval}(s,{\sf val})$ remains hidden from an adversary $\mathcal{A}$ that corrupts up to $t_{\sf sec}$ parties, where the precise notion of “hidden” depends on the application. For example, when $\sf TC$ is a threshold decryption scheme, i.e., ${\sf Eval}(s,{\sf val})$ outputs the decryption of the ciphertext ${\sf val}$ using secret key $s$, the Secrecy property requires that $\sf TC$ is semantically secure in the presence of an attacker $\mathcal{A}$ that corrupts up to $t$ parties.
[…]
Game 0.A.2 (Unpredictability Under Chosen Message Attack)
[…]
Definition 12 (Unpredictability Under Chosen Message Attack)
[…]
Appendix 0.B Blockchain-Native Threshold Cryptosystem
The following lemma is a direct implication of the blockchain-native threshold cryptosystem definition (Definition 8). Intuitively, a blockchain-native threshold cryptosystem should not reconstruct the $\sf TC$ output before a message is finalized.
Lemma 1
Proof
Appendix 0.C Tight Blockchain-Native Threshold Cryptosystem
0.C.1 Analysis of Algorithm 2
Theorem 0.C.1